how much money do u need to invest in two bank accounts one at 8% another at 10% so the gain the same interest in a year
Answer by mananth(16949)

the interest earned from both accounts is same
let one amount be x @ 8%
other be y @ 10%
0.08x=0.1y
x/y = 0.1/0.08
x/y = 10/8 ( multiply & divide by 100)
x/y = 5/4
The money invested will be in the ratio of 5:4
CHECK
10000 @ 8%= 800
8000@10%= 800